# Model Predictive Adversarial Imitation Learning for Planning from Observation

### v1

@@ D.2 Simulated Navigation Task Details
Figure 9: Visualization of Four Expert Trajectories in Navigation Task. Cars are initialized around (0,0)(0,0)( 0 , 0 ) and navigate towards (10,10)(10,10)( 10 , 10 ) where the “circling” behavior begins, as discussed in Section 4.1.
@@
OOD Experiment. OOD Energy in Figure 4 is computed as described by Liu et al. [35]. Namely, with respect to the expert data $d^{E}$, we fit a reference distribution using $\tilde{P}_{E}=\mathcal{N}(\bar{\mu}_{E},\bar{\Sigma}_{E})$. Then, the OOD energy is given by $E(s;p_{E})=\log p_{E}(s)$. Some limitations of this procedure can be observed given that ID points for GAIL do not receive as much reward as one might expect. However, this remains reasonable considering that the GAIL policy may forget ID behavior, which can be seen in Footnote 2 by agents clearly ID remaining static throughout the episode. Future work might better explore quantifying OOD towards measuring AIL generalization through direct usage of the discriminator.
@@ D.3 Predictive Model Learning Towards Generalizable MPAIL
D.4 De-aggregated Benchmark Results
Figure 10: Benchmark results from Figure 6 de-aggregated across number of demonstrations. MPAIL results for the Cartpole task is re-labeled and differentiated MPAIL (OM) to clarify that its Model is learned fully Online.
D.5 Ablations
@@
D.6 Towards High-Dimensional Tasks from Demonstration with Sampled-Based MPC
Figure 12: Isaac Lab Ant-v2 Experiment.
@@ D.6 Towards High-Dimensional Tasks from Demonstration with Sampled-Based MPC
“Remembering” locally optimal policies through a learned policy-like proposal distribution may help planning capabilities generalize to higher-dimensional spaces. Additionally, modeling dynamics in latent-space and using model ensembles have been shown to significantly improve performance in model-based reinforcement learning [49] and are promising directions for future work for high-dimensional tasks. Finally, Figure 8 illustrates a key takeaway of this framework: learning costs through MPAIL remains orthogonal to other works which seek to improve sample-based MPC through sampling [40, 56, 59], optimization [56, 60], and dynamics [20]. We believe that integration of developments in MPC along with application-specific cost regularization [51] may be critical for exploring the full potential of planning from observation.

### v2

@@ D.2 Simulated Navigation Task Details
OOD Experiment. OOD Energy in Figure 5 is computed as described by Liu et al. (2020). Namely, with respect to the expert data $d^{E}$, we fit a reference distribution using $\tilde{P}_{E}=\mathcal{N}(\bar{\mu}_{E},\bar{\Sigma}_{E})$. Then, the OOD energy is given by $E(s;p_{E})=\log p_{E}(s)$. Some limitations of this procedure can be observed given that ID points for GAIL do not receive as much reward as one might expect. However, this remains reasonable considering that the GAIL policy may forget ID behavior, which can be seen in Figure 4 by agents clearly ID remaining static throughout the episode. Future work might better explore quantifying OOD towards measuring AIL generalization through direct usage of the discriminator.
@@ D.3 Predictive Model Learning Towards Generalizable MPAIL
Figure 11: Benchmark results from Figure 7 de-aggregated across number of demonstrations. MPAIL results for the Cartpole task is differentiated MPAIL (OM) to clarify that its Model is learned fully Online.
D.4 Ablations
@@
D.5 Towards High-Dimensional Tasks from Demonstration with Sampled-Based MPC
Figure 13: Isaac Lab Ant-v2 Experiment with MPAIL (OM) and GAIL.
@@ D.5 Towards High-Dimensional Tasks from Demonstration with Sampled-Based MPC
“Remembering” locally optimal policies through a learned policy-like proposal distribution may help planning capabilities generalize to higher-dimensional spaces. Additionally, modeling dynamics in latent-space and using model ensembles have been shown to significantly improve performance in model-based reinforcement learning (Hansen et al., 2022) and are promising directions for future work for high-dimensional tasks. Finally, Figure 10 illustrates a key takeaway of this framework: learning costs through MPAIL remains orthogonal to other works which seek to improve sample-based MPC through sampling (Xue et al., 2024; Vlahov et al., 2024; Sacks and Boots, 2023), optimization (Vlahov et al., 2024; Sacks et al., 2024), and dynamics (Hansen et al., 2024). We believe that integration of developments in MPC along with application-specific cost regularization (Finn et al., 2016) may be critical for exploring the full potential of planning from observation.